# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ IV.2 Thermalization via instantons
Let us now address the flow away from the prethermal fixed point. For that matter, we study ${\left\|{H_{1}(\lambda)}\right\|}$ for different values of the next-nearest-neighbor coupling $J_{2}$, at and away from a freezing point, up to intermediate fRG time, $\lambda_{c}$. In this subsection, we will set $\lambda_{c}\gg\lambda_{\text{min}}$. We note that $J_{2}\neq 0$ breaks free-fermion integrability of the model. Therefore, we expect that $H_{1}(\lambda)$ ceases to decay, and shows an upturn at intermediate fRG times when $J_{2}\neq 0$. Our access to this regime is partly constrained by two factors: First, the initial exponential decay of $H_{1}(\lambda)$ is limited by numerical floating-point precision. The prethermal cutoff fRG time $\lambda_{\text{min}}$ and the drive frequency $\Omega$ are limited by $\Omega\lambda_{\text{min}}<\ln(1/\epsilon_{\text{fp}})$, where $\epsilon_{\text{fp}}\sim 10^{-16}$ is the floating-point precision. Beyond that, $H_{1}(\lambda)$ will decay below floating-point precision before reaching the upturn. Second, the growth of ${\left\|{H_{1}(\lambda)}\right\|}$ is associated with a non-trivial operator-spread in space, causing the bond dimension of the MPO to grow rapidly. Due to these constraints, we choose to study this regime using the ED method with a smaller value of $\Omega$, as shown in Fig. 6. The latter issue can be addressed using more sophisticated MPO compression in future work [85], but we expect that the qualitative considerations are likely to be similar to what we describe below.
@@
Clearly, as the system flows away from the prethermal fixed point, we expect the emergent symmetry associated with dynamical freezing to also be lost. However, we point out that dynamical freezing does provide additional mitigation against thermalization. As we see in Fig. 6, for a many-body system driven at the freezing point, the minimum fRG time $\lambda_{\text{min}}$ of ${\left\|{H_{1}(\lambda)}\right\|}$ is indeed delayed compared to the same system away from freezing. At the same time, ${\left\|{H_{1}(\lambda_{\text{min}})}\right\|}$ is smaller, and $H_{0}(\lambda)$ shows a longer plateau-like interval. This implies that the dynamically frozen regime (for the cosine-wave drive) thermalizes slower compared to a generic driven (unfrozen) system, in the sense of both fRG time and real time.
@@
As we will discuss in Sec. V, an instanton event is associated with the reorganization of the spectrum of $H_{0}(\lambda)$. When $\lambda\gg\lambda_{\text{min}}$, most instantons are elementary in the sense that they only involve a pair of eigenvalues of $H_{0}(\lambda)$ and a single matrix element of $H_{1}(\lambda)$, which is amenable to an analytical description. As a result, ${\left\|{H_{1}(\lambda)}\right\|}$ shows a simple ‘single-peak’ structure that agrees with the analytical solution in Eq. (22) below. However, we note there are also instances where multiple instantons are close to each other, and there can be nontrivial interactions between them. As shown in Fig. 7, the first growth regime of ${\left\|{H_{1}(\lambda)}\right\|}$ after $\lambda=\lambda_{\text{min}}$ is one of such examples, where two instantons are close to one another resulting in a ‘double-peak’ feature in ${\left\|{H_{1}(\lambda)}\right\|}$. In Appendix F, we present results for a smaller system size ($L=6$), where we were able to simulate the flow for significantly longer fRG times and access more instanton events. We found examples of well-isolated instantons, which continue to be well described by the theoretical results in Sec. V.
V Analytical results
@@
In this section, we present analytical results related to the solutions of the flow equations Eq. (3), especially in the putative dynamically frozen regime and beyond. Our discussion is divided into two parts. First, at early times, which includes the ramp-up regime and the approach to the prethermal fixed-point, the flow equation can be analyzed in a perturbative manner that is similar in spirit to the Magnus expansion, but with important key differences that appear in the beyond leading-order terms. Second, in the late time regime (which is the main regime of interest), the flow equation describes the departure from the prethermal fixed point and is analyzed non-perturbatively. This is where we obtain an approximate analytical handle on the instanton events that control the flow towards the thermal fixed-point.
V.1 Early time: approach to prethermal behavior
@@
The expansion in Eq. (13) clearly shares some similarity with the standard Floquet-Magnus expansion, which is believed to capture the asymptotic high-frequency behavior of a Floquet system [66]. For instance, the leading order term can be computed analytically, as discussed in detail in Appendix. C. Under the simplifying condition, $H_{1}(0)=H_{-1}(0)$, we obtain at leading order in $\zeta$:
$H_{0}^{(1)}(\lambda)$ $=$ $\left[J_{1}(\hat{z}e^{-\lambda\Omega})Y_{0}(\hat{z})-Y_{1}(\hat{z}e^{-\Omega\lambda})J_{0}(\hat{z})\right]$
@@
which serves as the leading order description of the prethermal fixed-point. Higher order corrections can be obtained order-by-order using the expansion in Eq. (13). However, we find that the higher order terms are different from the usual Magnus expansion. In particular, the correction term from the Magnus expansion for the present problem would be $\mathcal{O}(1/\Omega)$, but from the above flow expansion we obtain the correction to be $\mathcal{O}(1/\Omega^{2})$.
@@
Here $h_{m}$ is the Fourier coefficient of the co-moving frame Hamiltonian (see Appendix. C for more details). Note that Eq. (16) is explicitly gauge-dependent under a shift of $t_{0}\to t_{0}+\delta t$, where $h_{m}\to h_{m}\exp(im\Omega\delta t)$.
@@
Let us now apply the solution in Eq. (14) to explicitly address dynamical freezing. The setup is as follows: we choose the bare drive Hamiltonian, $H_{1}(0)=H_{-1}(0)=A{\mathcal{Q}}$. Here ${\mathcal{Q}}$ is the charge of the emergent symmetry, and $A$ is the driving amplitude, as before. We choose the bare static Hamiltonian to be, $H_{0}(0)=H_{0,{\mathcal{Q}}=0}+H_{0,{\mathcal{Q}}=1}$. As the subscripts suggest, the two parts have zero and unit eigenvalues, respectively, under the action of ${\mathcal{Q}}$: $[{\mathcal{Q}},H_{0,{\mathcal{Q}}=q}]=qH_{0,{\mathcal{Q}}=q}$. The static Hamiltonian associated with the prethermal fixed point is then,
@@
Here, we have chosen to flow the system to $\lambda=\lambda_{\text{pre}}$, such that $\Omega\lambda_{\text{pre}}\gg 1$, but but still short compared to the thermalization timescales, which will be discussed later. Therefore, by choosing $2A/\Omega=\nu_{i}$, where $\nu_{i}$ are zeros of the Bessel function $J_{0}(...)$, the prethermal Hamiltonian will display an (approximate) emergent symmetry $[H_{0}(\lambda_{\text{pre}}),{\mathcal{Q}}]=\mathcal{O}(1/\Omega^{2})$ in the high-frequency limit. To leading order, the locations of the freezing points agree with the prediction of the Magnus expansion.
The leading order solution in Eq. (14) also helps explains several features of the numerical results presented in Sec. IV. First, in the approach to the prethermal Hamiltonian, the commutator $[H_{0}(\lambda),{\mathcal{Q}}]$ shows dips whose number is given by $(n-1)$ for a freezing point, FPn, respectively. Substituting $\hat{z}=\nu_{i}$ into Eq. (14), we see that the dips are associated with zeros of $J_{1}(...)$, which correspond to extreme points of $J_{0}$, and between every zero of $J_{0}$ there is one extreme point. Second, to approach the prethermal fixed point, we require $(2A/\Omega)\exp(-\Omega\lambda)\ll 1$. Therefore, for higher freezing points corresponding to larger $A/\Omega$, it takes longer fRG time to reach the prethermal plateau.
However, as is well known, for a generic driven system that thermalizes to an infinite temperature state, the expansion in Eq. (13) must eventually break down; the higher-order terms in $H_{0}/\Omega$ should become increasingly relevant at a later fRG time. This connection between the breakdown of Eq. (13) and heating can be proven based on our earlier discussion of Eq. (5). Suppose there is no breakdown, such that the exponential decay of $H_{1}(\lambda)$ implies that $\exp(-itH_{0}(\lambda))$ is an increasingly better approximation of the actual time-evolution operator. However, this contradicts heating because $\exp(-itH_{0}(\lambda))$ conserves energy. In other words, the prethermal fixed point is unstable in a generic non-integrable system. In the next subsection, we present a qualitative and semi-analytic treatment of the eventual thermalization dynamics even at freezing, beyond the perturbative considerations of the present section.
V.2 Late-time: thermalization via instantons
@@
Expanding Eq. (18) in the eigenbasis $(E_{m},\ket{m})$ of $H_{0}(\lambda_{\text{pre}})$, we find
$\braket{m}{H_{1}(\lambda)}{n}\approx e^{(-\Omega-E_{m}+E_{n})(\lambda-\lambda_{\text{pre}})}\braket{m}{H_{1}(\lambda_{\text{pre}})}{n}\,.$ (19)
According to Eq. (19), $H_{1}(\lambda)$ will decay if $\Omega>|E_{m}-E_{n}|$ for all pairs of $\{m,n\}$ with $\braket{m}{H_{1}(\lambda_{\text{pre}})}{n}\neq 0$. This case corresponds to the pre-thermal fixed point being stable, and the associated expansion in Eq. (13) is expected to converge. However, this condition is not always satisfied, even for exponentially local Hamiltonians. Taking a spin chain as an example, an exponentially local drive term $H_{1}(\lambda_{\text{pre}})$ can contain terms that flip $l$ spins simultaneously for arbitrary $l$, despite its amplitude decaying exponentially in $l$. If the typical energy of a single spin flip is $J$, the matrix elements for $l>\mathcal{O}(\Omega/J)$ spin flips will eventually grow after a sufficiently long fRG time. In Appendix. D, we provide an estimate for these time scales associated with the evolution of ${\left\|{H_{1}(\lambda)}\right\|}$.
We argue that ${\left\|{H_{1}(\lambda)}\right\|}$ is expected to show a minimum at $\lambda=\lambda_{\text{min}}\sim(1/J)\ln(\Omega/J)$, where ${\left\|{H_{1}(\lambda)}\right\|}\sim\exp\left[-(\Omega/J)\ln(\Omega/J)\right]$; see Fig 9 in Appendix D for numerical evidence of such scaling. This implies that we can choose $\lambda_{\text{pre}}=\lambda_{\text{min}}$, and $H_{0}(\lambda_{\text{min}})$ is a good approximation of the time evolution up to exponentially long time in $\Omega$. This is consistent with generic expectations that heating in rapidly driven system is exponentially slow in $\Omega$ [30, 31, 32, 33, 34, 35, 36, 37, 38] . When $\lambda\gg\lambda_{\text{min}}$, we expect a growth in ${\left\|{H_{1}(\lambda)}\right\|}$, where different operators in $H_{1}(\lambda)$ are expected to have distinct thermalization timescales. For an operator that flips $l$ spins, the thermalization time scale is expected to be
@@
We can describe the evolution of $H_{1}(\lambda)$ approximately, under the assumption that only one of the matrix elements $\braket{m}{H_{1}(\lambda)}{n}$ grows to be $\mathcal{O}(1)$ — we dub this a single-instanton event. In this scenario, we ignore all of the other matrix elements of $H_{1}(\lambda)$. Then, the term $[H_{1},H_{1}^{\dagger}]$ in Eq. (3a) becomes approximately diagonal in the instantaneous eigenbasis of $H_{0}$. The flow equations simplify to
$\partial_{\lambda}E_{m}(\lambda)$ $=$ $2|\braket{m}{H_{1}(\lambda)}{n}|^{2}\,,$ (21a)
$\partial_{\lambda}E_{n}(\lambda)$ $=$ $-2|\braket{m}{H_{1}(\lambda)}{n}|^{2}\,,$ (21b)
$\partial_{\lambda}\braket{m}{H_{1}(\lambda)}{n}$ $=$ $[-\Omega-E_{m}(\lambda)+E_{n}(\lambda)]\braket{m}{H_{1}(\lambda)}{n}\,.$
Here $E_{m}(\lambda)$ and $E_{n}(\lambda)$ are the eigenvalues of the instanteous eigenvectors {|$\{\ket{m},\ket{n}\}$ of $H_{0}(\lambda)$, and |$\ket{m},\ket{n}$ are approximately independent of $\lambda$. The analytical solutions are given by,
$E_{n}(\lambda)-E_{m}(\lambda)$ $=$ $\Omega-\tilde{\Omega}\tanh(\tilde{\Omega}(\lambda-\tilde{\lambda}))\,,$ (22a)
$|\braket{m}{H_{1}(\lambda)}{n}|$ | $=$ $\frac{\tilde{\Omega}}{2\cosh(\tilde{\Omega}(\lambda-\tilde{\lambda}))}\,,$ (22b)
@@
For a generic non-integrable Hamiltonian, we expect the above single-instanton scenario to be a good approximation for the thermalization dynamics, in the absence of any special degeneracies in $E_{m}(\lambda)-E_{n}(\lambda)$, and when only one pair of levels are coupled strongly by $H_{1}(\lambda)$ via $\braket{m}{H_{1}(\lambda)}{n}$. Under this assumption, different matrix elements of $H_{1}(\lambda)$ will become ${\cal{O}}(1)$ at different fRG times, and at a given fRG time only one particular matrix element is large. As a corollary, near $\lambda=\tilde{\lambda}$, the norm ${\left\|{H_{1}(\lambda)}\right\|}$ will be dominated by a single matrix element. This line of reasoning and the associated analytical form of ${\left\|{H_{1}(\lambda)}\right\|}$ in Eq. (22b) is consistent with our numerical results, as shown by the excellent fit in Fig. 7 (Additional examples are also shown in Appendix F.). We note that in the presence of additional symmetries that enforce degeneracies in $E_{m}(\lambda)-E_{n}(\lambda)$, the Eqs. (21) can be extended to include the whole irreducible representation of the symmetry. We leave this generalization for future work.

### v2

@@ IV.3 Thermalization via instantons
Let us now address the flow away from the prethermal fixed point. For that matter, we study $\left\|{H_{1}(\lambda)}\right\|$} for different values of the next-nearest-neighbor coupling $J_{2}$, at and away from a freezing point, up to intermediate fRG time, $\lambda_{c}$. In this subsection, we will set $\lambda_{c}\gg\lambda_{\text{min}}$. We note that $J_{2}\neq 0$ breaks free-fermion integrability of the model. Therefore, we expect that $H_{1}(\lambda)$ ceases to decay, and shows an upturn at intermediate fRG times when $J_{2}\neq 0$. Our access to this regime is partly constrained by two factors: First, the initial exponential decay of $H_{1}(\lambda)$ is limited by numerical floating-point precision. The prethermal cutoff fRG time $\lambda_{\text{min}}$ and the drive frequency $\Omega$ are limited by $\Omega\lambda_{\text{min}}<\ln(1/\epsilon_{\text{fp}})$, where $\epsilon_{\text{fp}}\sim 10^{-16}$ is the floating-point precision. Beyond that, $H_{1}(\lambda)$ will decay below floating-point precision before reaching the upturn. Second, the growth of $\left\|{H_{1}(\lambda)}\right\|$} is associated with a non-trivial spreading of operators in space, causing the bond dimension of the MPO to grow rapidly. Due to these constraints, we choose to study this regime using the ED method with a smaller value of $\Omega$, as shown in Fig. 6. The latter issue can be addressed using more sophisticated MPO compression in future work [54], but we expect that the qualitative considerations are likely to be similar to what we describe below.
Figure 6: The Frobenius-norm of (a, b) $H_{1}$ as a function of fRG time for increasing $J_{2}$ for $L=10$. (a) Away from freezing, and for $J_{2}=0$, the system flows toward an integrable fixed point. For $J_{2}\neq 0$, at small fRG times the system first flows toward a prethermal fixed point, and eventually flows to a thermal fixed point. (b) At freezing, the system flows toward a prethermal fixed point for small and intermediate fRG times, controlled by the emergent approximate integrability, with the subsequent flow to the thermal fixed point delayed relative to (a). Other parameters are set to: $\Omega=2J$, $J=1$, $A/\Omega=0.3$ away from freezing, $A/\Omega=0.601$ at FP1.
@@
Clearly, as the system flows away from the prethermal fixed point, we expect the emergent symmetry associated with dynamical freezing to also be lost. However, we point out that dynamical freezing does provide additional mitigation against thermalization. As we see in Fig. 6, for a many-body system driven at the freezing point, the minimum fRG time $\lambda_{\text{min}}$ of $\left\|{H_{1}(\lambda)}\right\|$} is indeed delayed compared to the same system away from freezing. At the same time, $\left\|{H_{1}(\lambda_{\text{min}})}\right\|$} is smaller, and $H_{0}(\lambda)$ shows a longer plateau-like interval (See [1] for the corresponding figure of $H_{0}(\lambda)$). This implies that the dynamically frozen regime (for the cosine-wave drive) thermalizes slower compared to a generic driven (unfrozen) system, in the sense of both fRG time and real time.
Figure 7: (a) Frobenius norm of $H_{1}(\lambda)$ as a function of fRG time for $L=10$. Inset: isolated peak at late fRG time along with theoretical fit $(a/2){\rm sech}(a(\lambda-b))$]. Parameters: $J=1$, $J_{2}=0.2$, $\Omega=2.0$ at FP1. (b) Frobenius norm of $H_{0}(\lambda)$, with a step-like feature correlated with the peak in $H_{1}(\lambda)$ in (a). Additional results for $L=6$ are shown in Appendix C.
@@
As we will discuss in Sec. VI, an instanton event is associated with the reorganization of the spectrum of $H_{0}(\lambda)$. When $\lambda\gg\lambda_{\text{min}}$, most instantons are elementary in the sense that they only involve a pair of eigenvalues of $H_{0}(\lambda)$ and a single matrix element of $H_{1}(\lambda)$, which is amenable to an analytical description. As a result, $\left\|{H_{1}(\lambda)}\right\|$} shows a simple ‘single-peak’ structure that agrees with the analytical solution in Eq. (26) below. However, we note there are also instances where multiple instantons are close to each other, and there can be nontrivial interactions between them. As shown in Fig. 7, the first growth regime of $\left\|{H_{1}(\lambda)}\right\|$} after $\lambda=\lambda_{\text{min}}$ is one of such examples, where two instantons are close to one another resulting in a ‘double-peak’ feature in $\left\|{H_{1}(\lambda)}\right\|$}. In Appendix C, we present results for a smaller system size ($L=6$), where we were able to simulate the flow for significantly longer fRG times and access more instanton events. We found examples of well-isolated instantons, which continue to be well described by the theoretical results in Sec. VI. Finally, we remark that the analytical treatment in Sec. VI relates the width of the instanton with its height, but the exact location and the height of the instanton is determined by the microscopic details of the Hamiltonian. Nevertheless, we expect the first instanton to happen at a RG time scale $\lambda$ that is parametrically comparable to $\lambda_{\text{min}}$.
V Numerical diagnosis of thermalization
Our analysis thus far has established the emergence of prethermal regimes and instanton-mediated dynamics near dynamical freezing. However, we have not addressed the fundamentally important question of the extent to which the system thermalizes and loses memory of its initial state in the late-time limit. To address this question, we now present results for two complementary numerical diagnostics: (i) the Operator Entanglement Entropy (OPEE)  [93, 87] associated with $H_{0}(\lambda)$, the flowing effective Hamiltonian, which probes the build-up of operator non-locality and complexity, and (ii) the Diagonal Ensemble (DE) average of local observables, which captures the infinite-time limit of their expectation value after dephasing in the Floquet basis. The OPEE serves as a global measure that is sensitive to random-matrix-like behavior, while the DE average tests for the loss of memory of the initial state. Together, these quantities allow us to assess the extent of thermalization at and near freezing, and to identify regimes where prethermal behavior persists upto late times.
Figure 8: (a) The operator entanglement entropy (OPEE) for the operator $H_{0}(\lambda)$ is plotted as a function of RG time, both at and away from the freezing point ($L=10$, $\Omega/J=0.5$). Both of the curves approach the RMT value. (b) For $\Omega/J=1.0$, at the freezing point, the OPEE is significantly suppressed compared to the non-freezing point. (c) The saturation value of the OPEE is shown as a function of the driving frequency at the first freezing point, across different system sizes. As the frequency increases, the OPEE saturation value approaches sub-volume law. At lower frequencies, the OPEE scales as $\mathcal{O}(L)$. The RMT value is shown as dashed lines. Other parameters: $J=1$, $J_{2}=0.9$, $B_{x}=0.1$.
V.1 Operator entanglement entropy
[…]
$S[O]=-\sum_{i}\lambda_{i}^{2}\log\lambda_{i}^{2}.$ (15)
Our numerical results for OPEE are shown in Fig. 8. In panels (a) and (b), we show the evolution of the OPEE as a function of the RG time $\lambda$, for low $\Omega\lesssim\Omega_{*}$ and high frequency $\Omega\gtrsim\Omega_{*}$, where $\Omega_{*}$ is a system-size dependent “threshold” scale. In panel (c), we show the asymptotic saturation value of the OPEE as $\lambda\to\infty$ as a function of $\Omega$.
As shown in Fig. 8(a), when $\Omega\lesssim\Omega_{*}$ the OPEE appears to be insensitive to the difference between freezing and non-freezing points. Furthermore, the saturation value for OPEE agrees with that prescribed by random-matrix theory, with volume-law scaling [93]. On the other hand, for higher frequencies, the OPEE shows difference between freezing and non-freezing points as can be seen in Fig. 8(b). In particular, the OPEE at the freezing point saturates to a lower value compared to that at the non-freezing point. In Fig 8(c) we plot the saturation value of OPEE as a function of $\Omega$ across different system sizes. Moreover, this asymptotic saturation value deviates from the random-matrix prediction and exhibits a sub-volume-law scaling. This leads us to conjecture that the eventual thermalization behavior of the driven system, even at freezing, depends on how the drive frequency compares to a system-size dependent threshold, $\Omega_{*}(L)$.
To further investigate the thermalization properties in the thermodynamic limit $L\to\infty$, we need to analyze the $L$-dependence of $\Omega_{*}$. Based on the previous observations, if $\Omega_{*}(L)$ diverges with $L$ in the thermodynamic limit, it suggests that an infinite system driven at a finite frequency always thermalizes. On the other hand, if $\Omega_{*}(L\to\infty)$ is finite, it suggests that even an infinite system can avoid thermalization if driven at a sufficiently high but finite frequency. While our current numerical results are limited in size by ED, finite-size scaling suggests that $\Omega_{*}$ increases slowly with $L$, and a more systematic future analysis is required to settle this question definitively.
We note here that a quantum state at freezing can exhibit volume-law scaling of the state entanglement entropy [38], which has been attributed to Hilbert space fragmentation related to additional conserved quantities [36]. The OPEE that is studied here is a distinct concept from the state entanglement, which quantifies the spread of operator in the Liouville space. Saturation to the full RMT value we observed here for $\Omega<\Omega_{*}(L)$ is a significantly stronger diagnostic than the observation of a volume law alone: it requires unrestricted mixing across the entire operator Hilbert space, implying that $H_{0}(\lambda)$ explores the full Liouville space without constraint.
Our study of the OPEE has also revealed two additional features tied to freezing (see [1] for details), which include: (1) The OPEE for higher-order freezing points display an additional oscillatory behavior for short RG time $\lambda$ in the ramp-up regime, and the number of oscillations exactly matches the order of the freezing point minus one, which is reminiscent of the “dips” shown in Fig. 3(b). However, for later RG times the OPEE saturates to a value similar to the first freezing point, irrespective of the ratio of $\Omega/\Omega_{*}$. (2) The instanton events leave their fingerprints on the OPEE as “step” like features as a function of $\lambda$.
V.2 Diagonal ensemble average
To study how the memory of an initial state is preserved or lost for the driven Hamiltonian, we compute the diagonal ensemble average, which is intended to capture the expectation value of a local observable in the infinite-time limit. The DE average is defined as,
$\begin{split}\mathcal{O}_{\text{DEA}}&=\langle\Psi(NT\to\infty)|\mathcal{O}|\Psi(NT\to\infty)\rangle\\ &=\sum_{n}|c_{n}|^{2}\langle\phi_{n}|\mathcal{O}|\phi_{n}\rangle\end{split}$ (16)
[…]
Figure 9: Diagonal ensemble average of $S_{x}$ for a thermal pure state as a function of (a) driving frequency for different system sizes, (b) order of freezing for different driving frequencies. Other parameters: $L=10$, $B_{x}=0.1$, $J_{2}=0.9$.
VI Analytical results
@@
In this section, we present analytical results related to the solutions of the flow equations Eq. (3), especially in the putative dynamically frozen regime and beyond. Our discussion is divided into two parts. First, at early times, which includes the ramp-up regime and the approach to the prethermal fixed-point, the flow equation can be analyzed in a perturbative manner that is similar in spirit to the Floquet-Magnus expansion, but with important key differences that appear in the beyond leading-order terms. Second, in the late-time regime (which is the main regime of interest), the flow equation describes the departure from the prethermal fixed point and is analyzed non-perturbatively. This is where we obtain an approximate analytical handle on the instanton events that control the flow towards subsequent thermalization.
VI.1 Early time: approach to prethermal behavior
@@
The expansion in Eq. (17) clearly shares some similarity with the standard Floquet-Magnus expansion, which is believed to capture the asymptotic high-frequency behavior of a Floquet system [24]. For instance, the leading order term can be computed analytically, as discussed in detail in Appendix. D. Under the simplifying condition, $H_{1}(0)=H_{-1}(0)$, we obtain at leading order in $\zeta$:
[…]
@@
which serves as the leading order description of the prethermal fixed-point. Here, the $\lambda\to\infty$ limit can be refined to $)\lambda\gg(1/\Omega)\ln({\left\|{H_{1}(0)}\right\|}/(\Omega{\left\|{\hat{1}}\right\|}$)) based on Eq.(18). Higher order corrections can be obtained order-by-order using the expansion in Eq. (17). However, we find that the higher order terms are different from the usual Magnus expansion. In particular, the correction term from the Magnus expansion for the present problem would be $\mathcal{O}(1/\Omega)$, but from the above flow expansion we obtain the correction to be $\mathcal{O}(1/\Omega^{2})$.
@@
Here $h_{m}$ is the Fourier coefficient of the co-moving frame Hamiltonian (see Appendix. D for more details). Note that Eq. (20) is explicitly gauge-dependent under a shift of $t_{0}\to t_{0}+\delta t$, where $h_{m}\to h_{m}\exp(im\Omega\delta t)$.
@@
Let us now apply the solution in Eq. (18) to explicitly address dynamical freezing. The setup is as follows: we choose the bare drive Hamiltonian, $H_{1}(0)=H_{-1}(0)=A{\mathcal{Q}}$. Here $\mathcal{Q}$} is the charge of the emergent symmetry, and $A$ is the driving amplitude, as before. We choose the bare static Hamiltonian to be, $H_{0}(0)=H_{0,{\mathcal{Q}}=0}+H_{0,{\mathcal{Q}}=1}$. As the subscripts suggest, the two parts have zero and unit eigenvalues, respectively, under the action of $\mathcal{Q}$}: ${\mathcal{Q}},H_{0,{\mathcal{Q}}=q}]=qH_{0,{\mathcal{Q}}=q}$. The static Hamiltonian associated with the prethermal fixed point is then,
@@
[…]
However, as is well known, for a generic driven system that thermalizes to an infinite temperature state, the expansion in Eq. (17) must eventually break down; the higher-order terms in $H_{0}/\Omega$ should become increasingly relevant at a later fRG time. This connection between the breakdown of Eq. (17) and heating can be proven based on our earlier discussion of Eq. (5). Suppose there is no breakdown, such that the exponential decay of $H_{1}(\lambda)$ implies that $\exp(-itH_{0}(\lambda))$ is an increasingly better approximation of the actual time-evolution operator. However, this contradicts heating because $\exp(-itH_{0}(\lambda))$ conserves energy. In other words, the prethermal fixed point is unstable in a generic non-integrable system. In the next subsection, we present a qualitative and semi-analytic treatment of the subsequent dynamics even at freezing, beyond the perturbative considerations of the present section.
VI.2 Late-time: thermalization via instantons
@@
[…]
$\braket{m|H_{1}(\lambda)|n}\approx e^{(-\Omega-E_{m}+E_{n})(\lambda-\lambda_{\text{pre}})}\braket{m|H_{1}(\lambda_{\text{pre}})|n}\,.$ (23)
According to Eq. (23), $H_{1}(\lambda)$ will decay if $\Omega>|E_{m}-E_{n}|$ for all pairs of $\{m,n\}$ with $\braket{m|H_{1}(\lambda_{\text{pre}})|n}\neq 0$. This case corresponds to the pre-thermal fixed point being stable, and the associated expansion in Eq. (17) is expected to converge. However, this condition is not always satisfied, even for exponentially local Hamiltonians. Taking a spin chain as an example, an exponentially local drive term $H_{1}(\lambda_{\text{pre}})$ can contain terms that flip $l$ spins simultaneously for arbitrary $l$, despite its amplitude decaying exponentially in $l$. If the typical energy of a single spin flip is $J$, the matrix elements for $l>\mathcal{O}(\Omega/J)$ spin flips will eventually grow after a sufficiently long fRG time. In Appendix. E, we provide an estimate for these time scales associated with the evolution of $\left\|{H_{1}(\lambda)}\right\|$}.
We argue that $\left\|{H_{1}(\lambda)}\right\|$} is expected to show a minimum at $\lambda=\lambda_{\text{min}}\sim(1/J)\ln(\Omega/J)$, where ${\left\|{H_{1}(\lambda)}\right\|}\sim\exp\left[-(\Omega/J)\ln(\Omega/J)\right]$; see Fig 17 in Appendix E for numerical evidence of such scaling. This implies that we can choose $\lambda_{\text{pre}}=\lambda_{\text{min}}$, and $H_{0}(\lambda_{\text{min}})$ is a good approximation of the time evolution up to exponentially long time in $\Omega$. This is consistent with generic expectations that heating in rapidly driven system is exponentially slow in $\Omega$ [3, 46, 83, 40, 57, 72, 65, 56, 52] . When $\lambda\gg\lambda_{\text{min}}$, we expect a growth in $\left\|{H_{1}(\lambda)}\right\|$}, where different operators in $H_{1}(\lambda)$ are expected to have distinct thermalization timescales. For an operator that flips $l$ spins, the thermalization time scale is expected to be
@@
We can describe the evolution of $H_{1}(\lambda)$ approximately, under the assumption that only one of the matrix elements $\braket{m|H_{1}(\lambda)|n}$ grows to be $\mathcal{O}(1)$ — we dub this a single-instanton event. In this scenario, we ignore all of the other matrix elements of $H_{1}(\lambda)$. Then, the term $H_{1},H_{1}^{\dagger}]$ in Eq. (3a) becomes approximately diagonal in the instantaneous eigenbasis of $H_{0}$. The flow equations simplify to
$\partial_{\lambda}E_{m}(\lambda)$ $=$ $2|\braket{m|H_{1}(\lambda)|n}|^{2}\,$, (25a)
$\partial_{\lambda}E_{n}(\lambda)$ $=$ $-2|\braket{m|H_{1}(\lambda)|n}|^{2}\,$, (25b)
$\partial_{\lambda}\braket{m|H_{1}(\lambda)|n}$ $=$ $[-\Omega-E_{m}(\lambda)+E_{n}(\lambda)]\braket{m|H_{1}(\lambda)|n}\,.$
[…]
$E_{n}(\lambda)-E_{m}(\lambda)$ $=$ $),\displaystyle\Omega-\tilde{\Omega}\tanh(\tilde{\Omega}(\lambda-\tilde{\lambda}$))\,, (26a)
$|\braket{m|H_{1}(\lambda)|n}|$ $=$ $\frac{\tilde{\Omega}}{2\cosh(\tilde{\Omega}(\lambda-\tilde{\lambda}))}\,$, (26b)
@@
For a generic non-integrable Hamiltonian, we expect the above single-instanton scenario to be a good approximation for the thermalization dynamics, in the absence of any special degeneracies in $E_{m}(\lambda)-E_{n}(\lambda)$, and when only one pair of levels are coupled strongly by $H_{1}(\lambda)$ via $\braket{m|H_{1}(\lambda)|n}$. Under this assumption, different matrix elements of $H_{1}(\lambda)$ will become ${\cal{O}}(1)$ at different fRG times, and at a given fRG time only one particular matrix element is large. As a corollary, near $\lambda=\tilde{\lambda}$, the norm $\left\|{H_{1}(\lambda)}\right\|$} will be dominated by a single matrix element. This line of reasoning and the associated analytical form of $\left\|{H_{1}(\lambda)}\right\|$} in Eq. (26b) is consistent with our numerical results, as shown by the excellent fit in Fig. 7 (Additional examples are also shown in Appendix C.). We note that in the presence of additional symmetries that enforce degeneracies in $E_{m}(\lambda)-E_{n}(\lambda)$, the Eqs. (25) can be extended to include the whole irreducible representation of the symmetry. We leave this generalization for future work.